40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 3 16:43:44 2011, Katrin, Update, Green Locking, SHG temperature (YARM) 
    Reply  Thu Nov 3 17:04:23 2011, Koji, Update, Green Locking, SHG temperature (YARM) 
       Reply  Thu Nov 3 17:24:59 2011, Zach, Update, Green Locking, SHG temperature (YARM) 
          Reply  Fri Nov 4 13:53:33 2011, Katrin, Update, Green Locking, SHG temperature (YARM) 
    Reply  Wed Nov 9 16:49:17 2011, kiwamu, Update, Green Locking, Y end laser temperature with slow input connected 
Message ID: 5809     Entry time: Fri Nov 4 13:53:33 2011     In reply to: 5800
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: SHG temperature (YARM) 

Quote:

I must confess that I changed the temperature of the laser via the dial yesterday. I believe the initial (displayed) temperature was ~19o, whereas it is now probably in the high 20s. Sorry.

Quote:

Changing the crystal temperature changes the laser frequency. This will causes the beat note missing at the vertex.

In other words, you will find the beat note at the vertex when the actual temperature of the crystal is reproduced as before,
no matter how the dial setting/temp voltage input is.

 

Okay, my elog entry was not clear I changed the temperature of the SHG which only changes the conversion efficiency. 

 

 

Anyhow, since the laser set temperature and thus the laser frequency has been changed by Zach and I couldn't find a note

of the original laser crystal temperature, my plan is to reset the SHG temperature to the old value, set the laser crystal temperature

around 19°C and do fine adjustment of that temperature by optimising the doubling efficiency. Okay?

 

ELOG V3.1.3-